OpenStudy (loser66):

1)Compute \(e^{At}\) where \(A=\left[\begin{matrix}a&o\\b&c\end{matrix}\right]\) 2) Find the eigenvalues and eigenvectors of \(e^{-A}\) Please, help

OpenStudy (zzr0ck3r):

I forget this stuff. Do you need to diagonalize A first?

OpenStudy (loser66):

I am looking at my Discreet notes, now. Actually, it is from DE

OpenStudy (loser66):

hey barbecue, any idea??

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

I forget do i find reduced row echelon form for 1?

OpenStudy (loser66):

eigenvalues of A are a and c

OpenStudy (anonymous):

i don't know what I'm doing lol

OpenStudy (loser66):

eigenvectors are \(\left(\begin{matrix}a-c\\b\end{matrix}\right)\) and \(\left(\begin{matrix}0\\1\end{matrix}\right)\) Not sure about the second one, someone checks, please

OpenStudy (loser66):

Assume they are correct, then Diagonalization of A , namely \(D=\left[\begin{matrix}a&0\\0&c\end{matrix}\right]\) We get \(e^{At }= P*\left[\begin{matrix}e^{at}&0\\0&e^{ct}\end{matrix}\right]*P^{-1}\)

OpenStudy (loser66):

But I need confirm the eigenvalues before going further. @dan815 contribute, please

OpenStudy (loser66):

Then, we just do matrix multiplication to get the answer. ha!! but this knowledge is from Discreet, not from DE.